Jill is 5 years younger than his brother. 10 years from now, the sum of their ages will be 45, find Jill's current age
Answer by macston(5194)

.
Current sum of ages=45-20 years (10yrs for each)=25 years
J=Jill's age; B=brother's age=J+5yrs
.
J+B=25yrs Substitute for B.
J+J+5yrs=25 yrs
2J=20 yrs
J=10 yrs
ANSWER: Jill is 10 years old
.
CHECK:
Brother=J+5yrs=10yrs+5 yrs=15 yrs
Brother is 15 years old.
.
In 10 yrs:
(J+10yrs)+(B+10yrs)=45yrs
(10yrs+10yrs)+(15yrs+10yrs)=45yrs
20yrs+25yrs=45yrs
45yrs=45yrs